OpenStudy (anonymous):

Help me solve : 32/8 = 28/x

OpenStudy (mayankdevnani):

\[\huge \frac{32}{8}=\frac{28}{x}\] \[\huge 4=\frac{28}{x}\] by doing cross multiplying, \[\huge 4x=28\] dividing 4 from both the sides, \[\huge x=7\]